# Directed Acyclic Outerplanar Graphs Have Constant Stack Number

### v2

@@ Abstract
keywords:
Linear layouts, stack number, directed graphs, outerplanar graphs, 2-trees
@@ 1 Introduction
A directed acyclic graph, shortly DAG, is a directed graph with no directed cycles. For an integer $k\geq 1$ and a directed acyclic graph $G$, a $k$-stack layout of $G$ consists of a topological ordering $\prec$ of the vertices $V(G)$ and a partition of the edges $E(G)$ into $k$ parts such that each part is a stack. A part is a stack if no two of its edges cross with respect to $\prec$, where two edges $ab$ and $cd$ cross if their endpoints are ordered $a\prec c\prec b\prec d$ or $c\prec a\prec d\prec b$. Now the stack number111 The stack number is also known as book thickness or page number, especially in the older literature. In recent years the term stack number seems to be preferred over the others as it explicitly expresses the first-in-last-out property of every part, in alignment with related concepts like queue layouts. $\operatorname{sn}(G)$ of $G$ is the minimum $k$ such that there exists a $k$-stack layout of $G$. It is convenient to interpret the ordering $\prec$ as the vertices of $G$ being laid out from left to right, i.e., for $a\prec b$ we say that $a$ is to the left of $b$ (and $b$ is to the right of $a$) or that $a$ comes before $b$ in $\prec$ (and $b$ comes after $a$ in $\prec$). Then $\prec$ being a topological ordering of $G$ means that for every edge $ab$ directed from $a$ to $b$, vertex $a$ must come before vertex $b$ in $\prec$, in other words, every edge is directed from its left to its right endpoint according to $\prec$.
Given a topological ordering $\prec$ of $G$, a partition of $E(G)$ into $k$ stacks can equivalently be seen as a $k$-edge coloring, such that each color class is crossing-free. The simplest obstruction to admitting a partition into $k-1$ stacks is a set of $k$ pairwise crossing edges, also called a $k$-twist. For example, consider the $2k$-vertex graph $G_{k}$, $k\geq 2$, in the left of Figure 1 consisting of the directed path $P=(\ell_{1},\ldots,\ell_{k},r_{1},\ldots,r_{k})$ and the matching $M=\{\ell_{i}r_{i}\mid i=1,\ldots,k\}$. This directed acyclic graph has only one topological ordering $\prec$ in which the vertices are ordered along the directed path $P$. However with respect to this ordering $\prec$ the edges in $M$ form a $k$-twist. It follows that $\operatorname{sn}(G_{k})\geq k$ for all $k\geq 2$, which is also tight, as we can easily find a partition of $E(G_{k})$ into $k$ stacks as indicated in the right of Figure 1.
Figure 1: A planar directed acyclic graph $G_{k}$ of treewidth $3$ on $2k$ vertices (left) and its unique topological ordering $\prec$ containing a $k$-twist (right).
@@
For the sake of an example, let us argue that the outerplanar DAG $G$ in Figure 2 has $\operatorname{sn}(G)\geq 3$. In fact, due to symmetry it is enough to consider a topological ordering $\prec$ with $a_{3}$ to the left of $b_{1}$ and to observe that in $\prec$ the edges $a_{1}b_{1}$, $a_{2}b_{2}$, $a_{3}b_{3}$ form a $3$-twist. 1.1 concerning outerplanar DAGs, but also the more general question about directed acyclic $2$-trees, have received increasing attention over the last years (see Section 1.1 for more details), but it remained open to this day whether either of these classes contains DAGs of arbitrarily large stack number.
Figure 2: An outerplanar directed acyclic graph on eight vertices (left) and a topological ordering $\prec$ with $a_{3}$ to the left of $b_{1}$ containing a $3$-twist (right).
In addition to 1.1, there is a second major open question in the field of directed linear layouts which we attack. Here, a DAG is called upward planar if it can be drawn in the plane such that the edges are crossing-free and $y$-monotone.
Open Problem 1.2 (Nowakowski, Parker, 1989 [51]).
Is the stack number of the class of upward planar graphs bounded above by a constant?
@@
In this paper we answer both long-standing open problems connected to 1.1, that is whether or not outerplanar graphs and 2-trees have bounded stack number. We answer the first question in the positive, the second in the negative. In addition, we attack 1.2 and contribute a significant class of upward planar graphs with bounded stack number.
@@
In Section 3 we prove that 1.1 is true by showing that outerplanar DAGs have stack number at most $24776$. As one of our main tools we introduce directed $H$-partitions, which may be of independent interest for investigations of directed graphs in general. We remark that while another variant of $H$-partitions, so called layered $H$-partitions, caused a breakthrough in the investigation of queue layouts (a notion closely related to stack layouts) [24], this is – to the best of our knowledge – the first time that $H$-partitions are successfully adjusted to work with stack layouts.
Theorem 1.3.
@@
The stack number of outerplanar DAGs is bounded by a constant. Moreover, for every outerplanar DAG $G$ it holds that $\operatorname{sn}(G)\leq 24776$.
@@
Our second main result complements the constant upper bound for outerplanar DAGs by showing that already a just slightly larger, but still small, subclass of graphs of treewidth $2$ has unbounded stack number (see Section 2 for the formal definitions).
Theorem 1.4.
@@
The stack number of DAGs of treewidth $2$ is unbounded. Moreover, for every $k\geq 1$ there exists a monotone $2$-tree $G$ with $\operatorname{sn}(G)\geq k$ in which at most two vertices are stacked onto each edge.
@@
We remark that Theorems 1.3 and 1.4 together give a quite good idea for which 2-trees have bounded stack number as stacking at most one vertex onto each edge yields an outerplanar DAG. In $2006$, Di Giacomo, Didimo, Liotta and Wismath [23] asked whether or not all DAGs with treewidth 2 admit a $2$-stack layout. (The DAG in Figure 2 is already a counterexample.) Most recently, Nöllenburg and Pupyrev [52] highlight whether or not all directed acyclic $2$-trees have bounded stack number as an important open question. They as well as a Dagstuhl report by Bekos et al. [9] even conjecture that the stack number of all such $2$-trees should indeed be bounded. Our Theorem 1.4 refutes this conjecture.
@@
Let us also emphasize that Theorem 1.3 in particular gives that upward outerplanar graphs have bounded stack number. This was known only for specific subclasses before, such as internally triangulated upward outerpaths [17]. As such, Theorem 1.3 provides one of the largest known classes of upward planar graphs with bounded stack number, while it is a famous open problem whether or not actually all upward planar graphs have bounded stack number [51].
Organisation of the paper.
@@
Before proving Theorem 1.3 in Section 3 and Theorem 1.4 in Section 4, we define in Section 2 all concepts and notions relevant for our proofs. This includes a quick reminder of $2$-trees, outerplanar graphs, stack layouts, and twist numbers, but also some specialized notions for directed acyclic $2$-trees, such as monotone and transitive vertices, block-monotone DAGs, or transitive subgraphs below monotone vertices. Directed $H$-partitions are introduced in Section 3.2. But first, let us review related work.
@@ 1.1 Related Work
Stack layouts are just one type of so-called linear layouts which have been an active field of study over at least the last forty years. In full generality, a linear layout of an (undirected) graph $G$ consists of a total ordering $\prec$ of the vertices $V(G)$ and a partition of the edges $E(G)$ into parts such that each part fulfils certain combinatorial properties. For directed acyclic graphs the vertex ordering $\prec$ must be a topological ordering. The most prominent types of linear layouts are stack layouts (no two edges of the same part cross) and queue layouts (no two edges of the same part nest, i.e., two edges $ab$ and $cd$ ordered $a\prec c\prec d\prec b$ are forbidden) [40, 36, 24]. These are accompanied by a rich field of further variants like mixed layouts (a combination of stack and queue layouts) [40, 54, 6, 2, 21, 30], track layouts [25, 8, 55], deque layouts [7, 13, 15], rique layouts [11, 13, 15], local variants thereof [50, 49, 28, 6] and more [26, 1].
@@
[…]
@@
For directed acyclic graphs we additionally require the vertex ordering $\prec$ to be a topological ordering. Nowakowski and Parker [51] were the first to study this and consider stack layouts of diagrams of posets. Presenting an example with stack number $3$ (which was later improved to $4$ by Hung [41] and to $5$ by Merker [48]), they ask whether posets with a planar diagram [^3] In planar diagrams (generally: upward planar drawings) all edges must be drawn $y$-monotone along their edge direction. For example, the graph in Figure 2 is upward planar, while the graph $G_{k}$ in Figure 1 admits no upward planar drawing due to the edge from vertex $\ell_{k}$ to vertex $r_{1}$. have bounded stack number. Despite significant effort on different subclasses [57, 4, 37, 38, 39, 5, 3], this question still remains open.
@@
A slight generalization, namely whether all upward planar graphs have bounded stack number, is considered to be one of the most important open questions in the field of linear layouts [29, 52, 42]. It is known to hold for upward planar $3$-trees [29, 52]. However this does not imply the same for upward planar DAGs of treewidth at most $3$, since upward planar partial $3$-trees might not have an upward planar $3$-tree as a supergraph. (E.g., Figure 2 depicts such an example.) Superseding previous results [29], Jungeblut, Merker and Ueckerdt [42] recently gave the first sublinear upper bound for all upward planar graphs by showing that every $n$-vertex upward planar graph $G$ has stack number $\operatorname{sn}(G)\leq O((n\log n)^{2/3})=o(n)$.
@@
For general planar DAGs (that are not necessarily upward planar), Heath, Pemmaraju and Trenk [39] show that directed trees admit $1$-stack layouts and directed unicyclic graphs admit $2$-stack layouts. Other classes of DAGs admitting $2$-stack layouts include two-terminal series-parallel graphs [23], $N$-free graphs [47] or planar graphs whose faces have a special structure [18]. Recall from the example in Figure 1 that the stack number can be linear in the number of vertices, already for planar DAGs of treewidth $3$. With this in mind, Heath, Pemmaraju and Trenk [39] formulated 1.1 in $1999$.
@@
1.1 has received considerable attention, especially in recent years. Bhore, Da Lozzo, Montecchiani and Nöllenburg [17] confirm 1.1 for several subclasses of outerplanar DAGs, including internally triangulated upward outerpaths or cacti. Subsequently, Nöllenburg and Pupyrev [52] confirm 1.1 for single-source outerplanar DAGs, monotone outerplanar DAGs (to be defined in Section 2) and general outerpath DAGs. In a Dagstuhl Report [9], Bekos et al. claim that for every $n$-vertex outerplanar DAG $G$ it holds $\operatorname{sn}(G)\leq O(\log n)$, while they conjecture that actually $\operatorname{sn}(G)=O(1)$, even for all directed acyclic $2$-trees $G$. Theorem 1.3 in the present paper confirms 1.1, while Theorem 1.4 refutes the conjecture of Bekos et al.
[…]
@@ 2 Preliminaries
All graphs considered here are finite, non-empty, and have neither loops nor parallel edges.
@@
Outerplanar graphs are intimately related to graphs of treewidth $2$ and so-called $2$-trees. A $2$-tree is inductively defined through the following construction sequence:
  • A single edge $xy$ is a $2$-tree444 Usually this inductive definition of $2$-trees starts with a triangle, but for our methods it is more convenient to let a single edge be a $2$-tree as well. . This first edge in the process is called the base edge.
@@
Note that the construction sequence of a $2$-tree $G$ is not unique. In fact, for every $2$-tree $G$ and every edge $xy$ of $G$ there is a construction sequence of $G$ with $xy$ as the base edge. (And even for a fixed base edge, there can be exponentially many construction sequences of $G$.) However, as soon as the base edge $xy$ is fixed, this uniquely determines for each vertex $u$ different from $x$ and $y$ the edge $vw$ that $u$ is stacked onto. In this case we call $vw$ the parent edge of $u$, vertices $v$ and $w$ the parents of $u$, and likewise $u$ a child of $v$ and $w$. Note that for each edge $vw$ in $G$ different from the base edge, its endpoints $v$ and $w$ are in a parent/child relationship.
@@
Let $G$ be a connected outerplanar graph, see also Figure 2(a) for an example which is a subgraph of a simple $2$-tree (for now, ignore the edge orientations). A block $B$ of $G$ is either a bridge or a maximal $2$-connected component, while a vertex $v$ of $G$ is a cut vertex if $G-v$ is disconnected. The block-cut tree of $G$ has as vertex set all blocks and all cut vertices of $G$ and an edge between block $B$ and cut vertex $v$ if and only if $v\in B$. See Figure 2(b) for a block-cut tree of the graph in Figure 2(a).
@@
Figure 3: An outerplanar DAG $G$ with its block-cut tree. The construction sequence of $G$ is $1,2,\ldots,8$. In particular, the base edge is from $1$ to $2$. Dotted edges are non-edges of $G$ whose addition to $G$ gives an outerplanar $2$-tree.
@@
Observe that the notions of cyclic, transitive and monotone vertices crucially depend on the choice of the base edge. Further observe that a directed $2$-tree $G$ is a DAG if and only if no vertex is cyclic; independent of the choice of the base edge. So if $G$ is a DAG, then every construction sequence involves only transitive and monotone vertices.
We say that a directed $2$-tree $G$ is monotone (respectively transitive) if there exists a choice for the base edge $xy$ such that every vertex except for $x$ and $y$ is monotone (respectively transitive). In particular, a monotone (or transitive) directed outerplanar graph is always a maximal directed outerplanar graph. A connected outerplanar (but not necessarily maximal outerplanar) DAG is called block-monotone if every block is monotone, see e.g., Figure 5.
Figure 5: A block-monotone outerplanar DAG $G$. Each of the two blocks of $G$ contains a base edge ($xy$ and $uv$) such that all other vertices are monotone.
Given a maximal outerplanar DAG $G$ with a fixed base edge $xy$, its construction tree with respect to $xy$ is a rooted, undirected (and unordered) binary tree $T$ on the vertices of $G$ with vertex labels M (for ?monotone?) and T (for ?transitive?) such that555 We remark that the construction tree (unrooted and without the labels) is exactly the tree of a nice tree-decomposition of width $2$ of $G$, but we do not use this fact here. :
@@
Figure 6: An outerplanar graph and its construction tree. The transitive subgraph below each monotone vertex is highlighted orange.
Stack numbers versus twist number.
@@
In fact, Davies [20] gives an upper bound on the chromatic number $\chi(H)$ of a circle graph $H$ in terms of its clique number $\omega(H)$. (Gyárfás [33] was the first to show that circle graphs are $\chi$-bounded but Davies gives the first asymptotically tight bound.) To obtain Theorem 2.1, simply consider the circle graph $H$ with $V(H)=E(G)$ whose edges correspond to crossing edges in $\prec$. Then $\omega(H)$ is the twist number of $\prec$ and a proper $k$-coloring of $H$ is a partition of $E(G)$ into $k$ stacks with respect to $\prec$.

### v3

@@ 1 Introduction
A directed acyclic graph, shortly DAG, is a directed graph with no directed cycles. For an integer $k\geq 1$ and a directed acyclic graph $G$, a $k$-stack layout of $G$ consists of a topological ordering $\prec$ of the vertices $V(G)$ and a partition of the edges $E(G)$ into $k$ parts such that each part is a stack. A part is a stack if no two of its edges cross with respect to $\prec$, where two edges $ab$ and $cd$ cross if their endpoints are ordered $a\prec c\prec b\prec d$ or $c\prec a\prec d\prec b$. Now the stack number [^1] The stack number is also known as book thickness or page number, especially in the older literature (stacks are called pages then). In recent years the term stack number seems to be preferred over the others as it explicitly expresses the first-in-last-out property of every part, in alignment with related concepts like queue layouts. $\operatorname{sn}(G)$ of $G$ is the minimum $k$ such that there exists a $k$-stack layout of $G$. It is convenient to interpret the ordering $\prec$ as the vertices of $G$ being laid out from left to right, i.e., for $a\prec b$ we say that $a$ is to the left of $b$ (and $b$ is to the right of $a$) or that $a$ comes before $b$ in $\prec$ (and $b$ comes after $a$ in $\prec$). Then $\prec$ being a topological ordering of $G$ means that for every edge $ab$ directed from $a$ to $b$, vertex $a$ must come before vertex $b$ in $\prec$, in other words, every edge is directed from its left to its right endpoint according to $\prec$.
Given a topological ordering $\prec$ of $G$, a partition of $E(G)$ into $k$ stacks can equivalently be seen as a $k$-edge coloring, such that each color class is crossing-free. The simplest obstruction to admitting a partition into k1k-1 stacks is a set of $k$ pairwise crossing edges, also called a $k$-twist. For example, consider the $2k$-vertex graph $G_{k}$, $k\geq 2$, in the left of Section 1 consisting of the directed path $P=(\ell_{1},\ldots,\ell_{k},r_{1},\ldots,r_{k})$ and the matching $M=\{\ell_{i}r_{i}\mid i=1,\ldots,k\}$. This directed acyclic graph has only one topological ordering $\prec$ in which the vertices are ordered along the directed path $P$. However with respect to this ordering $\prec$ the edges in $M$ form a $k$-twist. It follows that $\operatorname{sn}(G_{k})\geq k$ for all $k\geq 2$, which is also tight, as we can easily find a partition of $E(G_{k})$ into $k$ stacks as indicated in the right of Section 1.
{sidefigure}[Uncaptioned image][Uncaptioned image]
A planar directed acyclic graph $G_{k}$ of treewidth $3$ on $2k$ vertices (left) and its unique topological ordering $\prec$ containing a $k$-twist (right).
@@
For the sake of an example, let us argue that the outerplanar DAG $G$ in Conjecture 1.1 has $\operatorname{sn}(G)\geq 3$. In fact, due to symmetry it is enough to consider a topological ordering $\prec$ with $a_{3}$ to the left of $b_{1}$ and to observe that in $\prec$ the edges $a_{1}b_{1}$, $a_{2}b_{2}$, $a_{3}b_{3}$ form a $3$-twist. Conjecture 1.1 concerning outerplanar DAGs, but also the more general question about directed acyclic $2$-trees, have received increasing attention over the last years (see Section 1.1 for more details), but it remained open to this day whether either of these classes contains DAGs of arbitrarily large stack number.
{sidefigure}[Uncaptioned image][Uncaptioned image]
An outerplanar directed acyclic graph on eight vertices (left) and a topological ordering $\prec$ with $a_{3}$ to the left of $b_{1}$ containing a $3$-twist (right).
In addition to Conjecture 1.1, there is a second major open question in the field of directed linear layouts which we attack. Here, a DAG is called upward planar if it can be drawn in the plane such that the edges are crossing-free and $y$-monotone.
{open}
[Nowakowski, Parker, 1989 [52]] Is the stack number of the class of upward planar graphs bounded above by a constant?
@@
In this paper, we answer both long-standing open problems connected to Conjecture 1.1, that is whether or not outerplanar graphs and 2-trees have bounded stack number. We answer the first question in the positive, the second in the negative. Additionally, we attack Conjecture 1.1 and contribute a significant class of upward planar graphs with bounded stack number.
@@
In Section 3, we prove that Conjecture 1.1 is true by showing that outerplanar DAGs have stack number at most $24776$. As one of our main tools we introduce directed $H$-partitions, which may be of independent interest for investigations of directed graphs in general. We remark that while another variant of $H$-partitions, so called layered $H$-partitions, caused a breakthrough in the investigation of queue layouts (a notion closely related to stack layouts) [25], this is – to the best of our knowledge – the first time that $H$-partitions are successfully adjusted to work with stack layouts.
{restatable}
@@
theoremouterplanarBounded The stack number of outerplanar DAGs is bounded by a constant. Moreover, every outerplanar DAG $G$ has $\operatorname{sn}(G)\leq 24776$.
@@
Our second main result complements the constant upper bound for outerplanar DAGs by showing that already a just slightly larger subclass of graphs of treewidth $2$ has unbounded stack number. [^3] We remark that every 2-tree can be constructed by iteratively stacking vertices onto an edge, i.e., by introducing a new vertex that is connected exactly to the endpoints of an edge. A directed 2-tree is called monotone if in each step the new vertex is a source or a sink. See also Section 2.
{restatable}
@@
theoremtwotreesUnbounded The stack number of DAGs of treewidth $2$ is unbounded. Moreover, for every $k\geq 1$ there exists a monotone $2$-tree $G$ with $\operatorname{sn}(G)\geq k$ in which at most two vertices are stacked onto each edge.
@@
We remark that Sections 1 and 1 together give a quite good understanding of which 2-trees have bounded stack number as stacking at most one vertex onto each edge yields exactly the maximal outerplanar DAGs. In $2006$, Di Giacomo, Didimo, Liotta and Wismath [23] asked whether or not all DAGs with treewidth 2 admit a $2$-stack layout. (The DAG in Conjecture 1.1 is already a counterexample.) ¡Most recently, Nöllenburg and Pupyrev [51] highlight whether or not all directed acyclic $2$-trees have bounded stack number as an important open question. They as well as a Dagstuhl report by Bekos et al. [9] even conjecture that the stack number of all such $2$-trees should indeed be bounded. Our Section 1 refutes this conjecture.
@@
Let us also emphasize that Section 1 in particular gives that upward outerplanar graphs have bounded stack number. This was known only for specific subclasses before, such as internally triangulated upward outerpaths [17]. As such, Section 1 provides one of the largest known classes of upward planar graphs with bounded stack number, while it is a famous open problem whether or not actually all upward planar graphs have bounded stack number [52].
Organization of the paper.
@@
Before proving Section 1 in Section 3 and Section 1 in Section 4, we define in Section 2 all concepts and notions relevant for our proofs. This includes a quick reminder of $2$-trees, outerplanar graphs, stack layouts, and twist numbers, but also some specialized notions for directed acyclic $2$-trees, such as monotone and transitive vertices, block-monotone DAGs, or transitive subgraphs below monotone vertices. Directed $H$-partitions are introduced in Section 3.2. But first, let us review related work.
@@ 1.1 Related Work
Stack layouts are just one type of so-called linear layouts which have been an active field of study over at least the last forty years. In full generality, a linear layout of an (undirected) graph $G$ consists of a total ordering $\prec$ of the vertices $V(G)$ and a partition of the edges $E(G)$ into parts such that each part fulfills certain combinatorial properties. For directed acyclic graphs the vertex ordering $\prec$ must be a topological ordering. The most prominent types of linear layouts are stack layouts (no two edges of the same part cross) and queue layouts (no two edges of the same part nest, i.e., two edges $ab$ and $cd$ ordered $a\prec c\prec d\prec b$ are forbidden) [40, 36, 25]. These are accompanied by a rich field of further variants like mixed layouts (a combination of stack and queue layouts) [40, 55, 6, 2, 21, 29], track layouts [26, 8, 54], deque layouts [7, 13, 15], rique layouts [11, 13, 15], local variants thereof [50, 49, 28, 6] and more [27, 1].
@@
It is well-known that a graph with at least one edge admits a $1$-stack layout if and only if it is outerplanar, and that it admits a $2$-stack layout if and only if it is planar sub-Hamiltonian (i.e., is a subgraph of a planar graph containing a Hamiltonian cycle) [16]. Testing whether a graph is planar sub-Hamiltonian is $\mathsf{NP}$-complete [61], but there has been significant effort to identify planar Hamiltonian and sub-Hamiltonian graph classes: These include (among others) planar bipartite graphs [22], planar graphs with maximum degree at most $4$ [12], planar $4$-connected graphs [59], planar $3$-connected graphs with maximum degree $5$, and $2$-trees [56]. Further, three stacks are sufficient for planar $3$-trees [35] and planar graphs with maximum degree $5$ [32]. Recall that four stacks are always sufficient [63] and sometimes necessary [14, 64] for all planar graphs.
@@
For directed acyclic graphs we additionally require the vertex ordering $\prec$ to be a topological ordering. Nowakowski and Parker [52] were the first to study this and consider stack layouts of diagrams of posets. Presenting an example with stack number $3$ (which was later improved to $4$ by Hung [41] and to $5$ by Merker [48]), they ask whether posets with a planar diagram [^4] In planar diagrams (generally: upward planar drawings) all edges must be drawn $y$-monotone along their edge direction. For example, the graph in Conjecture 1.1 is upward planar, while the graph $G_{k}$ in Section 1 admits no upward planar drawing due to the edge from vertex $\ell_{k}$ to vertex $r_{1}$. have bounded stack number. Despite significant effort on different subclasses [57, 4, 38, 37, 39, 5, 3], this question still remains open.
@@
A slight generalization, namely whether all upward planar graphs have bounded stack number, is considered to be one of the most important open questions in the field of linear layouts [30, 51, 42]. It is known to hold for upward planar $3$-trees [30, 51]. However, this does not imply the same for upward planar DAGs of treewidth at most $3$, since upward planar partial $3$-trees might not have an upward planar $3$-tree as a supergraph. (E.g., Conjecture 1.1 depicts such an example.) Superseding previous results [30], Jungeblut, Merker and Ueckerdt [42] recently gave the first sublinear upper bound for all upward planar graphs by showing that every $n$-vertex upward planar graph $G$ has stack number $\operatorname{sn}(G)\leq O((n\log n)^{2/3})=o(n)$.
@@
For general planar DAGs (that are not necessarily upward planar), Heath, Pemmaraju and Trenk [39] show that directed trees admit $1$-stack layouts and directed unicyclic graphs admit $2$-stack layouts. Other classes of DAGs admitting $2$-stack layouts include two-terminal series-parallel graphs [23], $N$-free graphs [47] or planar graphs whose faces have a special structure [18]. Recall from the example in Section 1 that the stack number can be linear in the number of vertices, already for planar DAGs of treewidth $3$. With this in mind, Heath, Pemmaraju and Trenk [39] formulated Conjecture 1.1 in $1999$.
@@
Conjecture 1.1 has received considerable attention, especially in recent years. Bhore, Da Lozzo, Montecchiani and Nöllenburg [17] confirm Conjecture 1.1 for several subclasses of outerplanar DAGs, including internally triangulated upward outerpaths or cacti. Subsequently, Nöllenburg and Pupyrev [51] confirm Conjecture 1.1 for single-source outerplanar DAGs, monotone outerplanar DAGs (to be defined in Section 2) and general outerpath DAGs. In a Dagstuhl Report [9], Bekos et al. claim that every $n$-vertex outerplanar DAG $G$ has $\operatorname{sn}(G)\leq O(\log n)$, while they conjecture that actually $\operatorname{sn}(G)=O(1)$, even for all directed acyclic $2$-trees $G$. Section 1 in the present paper confirms Conjecture 1.1, while Section 1 refutes the conjecture of Bekos et al.
Finally, let us mention that the decision problem of whether a given DAG admits a $k$-stack layout is known to be $\mathsf{NP}$-complete for every $k\geq 2$ [10, 18, 37]. However, there are FPT algorithms parameterized by the branchwidth [18] or the vertex cover number [17].
@@ 2 Preliminaries
All graphs considered here are finite, non-empty, and simple, i.e., have neither loops nor parallel edges. For a graph $G$ and an edge $e$, we define $G+e$, respectively GeG-e, as the graph obtained from $G$ by adding, respectively removing, the edge $e$ if possible. Similarly, vertices, sets of edges or vertices, and subgraphs can be added or removed with the same notation, where incident edges are also removed as necessary.
@@
Outerplanar graphs are intimately related to graphs of treewidth $2$ and so-called $2$-trees. A $2$-tree is inductively defined by the following construction sequence:
  • A single edge $xy$ is a $2$-tree [^5] Usually this inductive definition of $2$-trees starts with a triangle, but for the arguments below it is more convenient to let a single edge be a $2$-tree as well. . This first edge in the process is called the base edge.
@@
We remark that 2-trees are exactly the edge-maximal graphs of treewidth 2. Note that the construction sequence of a $2$-tree $G$ is not unique. In fact, for every $2$-tree $G$ and every edge $xy$ of $G$ there is a construction sequence of $G$ with $xy$ as the base edge. (And even for a fixed base edge, there can be exponentially many construction sequences of $G$.) However, as soon as the base edge $xy$ is fixed, this uniquely determines for each vertex $u$ different from $x$ and $y$ the edge $vw$ that $u$ is stacked onto. In this case we call $vw$ the parent edge of $u$, vertices $v$ and $w$ the parents of $u$, and likewise $u$ a child of $v$ and $w$. Note that for each edge $vw$ in $G$ different from the base edge, its endpoints $v$ and $w$ are in a parent/child relationship.
@@
Let $G$ be a connected outerplanar graph, see also Figure 1(a) for an example which is a subgraph of a simple $2$-tree (for now, ignore the edge orientations). A block $B$ of $G$ is either a bridge or a maximal $2$-connected component, while a vertex $v$ of $G$ is a cut vertex if GvG-v is disconnected. The block-cut tree of $G$ has as vertex set all blocks and all cut vertices of $G$ and an edge between block $B$ and cut vertex $v$ if and only if $v\in B$. See Figure 1(b) for a block-cut tree of the graph in Figure 1(a).
@@
Figure 1: An outerplanar DAG $G$ with its block-cut tree. The construction sequence of $G$ is $1,2,\ldots,8$. In particular, the base edge is from $1$ to $2$. Dotted edges are non-edges of $G$ whose addition to $G$ gives an outerplanar 22-tree.
@@
Observe that the notions of cyclic, transitive and monotone vertices crucially depend on the choice of the base edge. Further, observe that a directed $2$-tree $G$ is a DAG if and only if no vertex is cyclic, independent of the choice of the base edge. So if $G$ is a DAG, then every construction sequence involves only transitive and monotone vertices.
We say that a directed $2$-tree $G$ is monotone (respectively transitive) if there exists a choice for the base edge $xy$ such that every vertex except for $x$ and $y$ is monotone (respectively transitive). In particular, a monotone (or transitive) directed outerplanar graph is always a maximal directed outerplanar graph. A connected outerplanar (but not necessarily maximal outerplanar) DAG is called block-monotone if every block is monotone, see e.g., Figure 2. {csidefigure} [Uncaptioned image] A block-monotone outerplanar DAG $G$. Each of the two blocks of $G$ contains a base edge ($xy$ and $uv$) such that all other vertices are monotone.
Given a maximal outerplanar DAG $G$ with a fixed base edge $xy$, its construction tree with respect to $xy$ is a rooted, undirected (and unordered) binary tree $T$ on the vertices of $G$ with vertex labels M (for “monotone”) and T (for “transitive”) such that [^6] We remark that the construction tree (unrooted and without the labels) is exactly the tree of a nice tree-decomposition of width $2$ of $G$, but we do not use this fact here. :
@@
{csidefigure}[Uncaptioned image]
An outerplanar graph with base edge $ab$ and its construction tree. The transitive subgraph below each monotone vertex is highlighted orange.
Stack number versus twist number.
@@
In fact, Davies [20] gives an upper bound on the chromatic number $\chi(H)$ of a circle graph $H$ in terms of its clique number $\omega(H)$. (Gyárfás [33] was the first to show that circle graphs are $\chi$-bounded, but Davies gives the first asymptotically tight bound.) To obtain Theorem 2.1, simply consider the circle graph $H$ with $V(H)=E(G)$ whose edges correspond to crossing edges in $\prec$. Then $\omega(H)$ is the twist number of $\prec$ and a proper $k$-coloring of $H$ is a partition of $E(G)$ into $k$ stacks with respect to $\prec$.